Suppose that the supply curve for econometrists is given by w = 10 + 5L, while the demand curve is given by w = 50 - 3L, where w = annual earnings in thousands of dollars per year and L = thousands of econometrists.

A. Find the equilibrium wage and employment levels.
B. Now suppose that the demand for econometrists increases and the new demand curve is w = 70 – 3L. Assume that this market is subject to cobwebs because it takes about three years to produce people who specialize in econometrics. While this adjustment is taking place, the short –run supply of econometrists is fixed. Calculate the wage and employment levels in each of the first five rounds.
C. Find the new long-run equilibrium wage and employment levels
D. Draw a graph of the cob web model above with clear labeling of wage and corresponding employment level in each round.
1
Expert's answer
2019-04-18T09:28:51-0400

A. Employment levels are equal at equilibrium wage, so:

10 + 5L = 50 - 3L,

8L = 40,

L = 5 workers,

w = 10 + 5*5 = 35 thousands of dollars.

B. The wage and employment levels in each of the first five rounds can be found from the new demand curve.

C. If the new demand curve is w = 70 – 3L, then the new equilibrium wage and employment levels are:

10 + 5L = 70 - 3L,

8L = 80,

L = 10 workers,

w = 10 + 5*10 = 60 thousands of dollars.

D. The cob web model is an economic model that explains why prices might be subject to periodic fluctuations in certain types of markets. A graph will show the demand and supply curves, which will intersect in equilibrium point.